f(x)=-2x^3+5 find x if f)x)=-49
Answer by Cromlix(4381)

-2x^3 + 5 = -49
-2x^3 = -49 -5
-2x^3 = -54
2x^3 = 54 (Multiply thro' by -1)
x^3 = 27
x = 3
Hope this helps.
:-)